OpenStudy (anonymous):

Please help with these two functions, I have to solve for x -2=log(3x+5) and e^x e^(x+1)=1 Thank you!!!

OpenStudy (nincompoop):

you're welcome

OpenStudy (shamil98):

Use the rule: \[\huge \log_a x = n ---> x = a^n\] Treat (3x+5) as a variable. if there is no base log there, we assume it has the base of 10. so \[\huge \log_{10}(3x+5) = -2\] Use the given rule above and do it from there :)

OpenStudy (mathmale):

Shami is on the right track. The principle here is that y = (log to the base a of x) and y = a^x are INVERSE FUNCTIONS. Thus, 3 raised to the power (log to the base a of x) results in just x. Next, 3 raised to the power -2 results in what numerical value? Write the solution: x = 3^(-2). Apply the same procedure to solve Emmm's original question.